OpenStudy (anonymous):

f(x)=f'(x) find f(x) i knw 2 f(x)=e^x and f(x)=0 can u tell me more

OpenStudy (anonymous):

so you are looking for functions that equal the derivative of the function? f(x)=e^x, f'(x)=e^x, f(x)=0, f'(x)=0 both like you said. What about undefined functions? Can those be used? f(x)=x/0 = infinity, and the derivative would also equal infinity. That's probably not very good.

OpenStudy (anonymous):

no i dont prefer undefined functions and i dont want functions like this either f(x)=x - x as this is another form of f(x)=0